Use the following equation: y = –3(x + 8)^2 + 2
The coordinates of the vertex written as (h, k) would be?

Answer by Boreal(15235)

They are (-8,2). You reverse the sign of the constant in the parentheses and take the last constant without changing sign.
This is also
f(x)=-3x^2-48x-190.
The vertex is -b/2a which is 48/-6 =-8
f(x)=-192+384-190=2